# Patch-based learning of adaptive Total Variation parameter maps for blind image denoising

### v1

@@ Abstract
We consider a patch-based learning approach defined in terms of neural networks to estimate spatially adaptive regularisation parameter maps for image denoising with weighted Total Variation and test it to situations when the noise distribution is unknown. As an example, we consider situations where noise could be either Gaussian or Poisson and perform preliminary model selection by a standard binary classification network. Then, we define a patch-based approach where at each image pixel an optimal weighting between TV regularisation and the corresponding data fidelity is learned in a supervised way using reference natural image patches upon optimisation of SSIM and in a sliding window fashion. Extensive numerical results are reported for both noise models, showing significant improvement w.r.t. results obtained by means of optimal scalar regularisation.
@@ I Introduction
More recent approaches (see, e.g., [7]) consider weighted, that is spatially-adaptive, variants of TV-based models where a pixel-dependent weighting between regularisation and data term is used to account for differences between textured and geometric features, thus avoiding unbalanced effects corresponding to a global selection of $\mu$, see Fig. 1, where the regularisation parameter maps is computed in a patch-wise fashion and optimal weighting values are computed on an exemplar image corrupted by Gaussian noise with $\sigma^{2}=0.01$.
@@ I-A Contribution
We consider a patch-based learning approach for estimating adaptive, i.e. spatially-varying regularisation maps for the problem of image denoising with unknown noise distribution. The regularisation map $\bm{\mu}_{\text{map}}$ is estimated in a supervised learning fashion where, for each noisy patch $\bm{y}_{p}$, the map $\bm{y}_{p}\mapsto\mu_{i}$ providing an estimate of the regularisation parameter to-be-located in the pixel $i=1,\ldots,n$ placed at the centre of $\bm{y}_{p}$ is parametrised by a neural $f_{\theta}$ network which, at inference time, operates in a sliding-window fashion, taking as inputs patches of fixed size. An example of the proposed map estimation can be appreciated in Fig. 2.
Figure 2: Pipeline of our proposed method.
Noise model selection is performed using a standard binary classification network able to distinguish between Gaussian and Poisson noise. Several numerical results are reported showing visual and statistical improvements w.r.t. non-adaptive regularisation models.
@@ II Patch-based learning of regularisation maps
We consider in the following an adaptive approach that assigns to each pixel $i\in\left\{1,\ldots n\right\}$ of a noisy image $\bm{y}$ corrupted by Gaussian/Poisson noise a scalar regularisation parameter $\mu_{i}$ computed by means of a suitable training of a deep neural network leveraging information on a small patch centered in the pixel itself. More precisely, given a patch $\bm{y}_{p_{i}}$ extracted by $\bm{y}$ around $i\in\left\{1,\ldots n\right\}$, we consider the following patch-wise version of problem (3) suited to deal with Gaussian/Poisson which reads:
@@
where $\delta\in\left\{0,1\right\}$ identifies (upon suitable training) whether Gaussian/Poisson noise is observed in the image, so that a tailored data term is used for reconstruction. The parameter $\mu_{i}$ is the one we would like to estimate to construct the desired regularisation map. We now parametrise the map $\bm{y}_{p_{i}}\mapsto\mu_{i}$ using a deep neural network $f_{\bm{\theta}}$. In the following, we describe the supervised-learning strategy which, given a dataset of clean/noisy natural image patches $\left\{(\tilde{\bm{x}}_{p}^{k},\bm{y}^{k}_{p})\right\}_{k=1}^{K}$ learn such map. At inference time, note that the this map will be applied in a sliding-window fashion, see Fig. 3, to an unobserved image $\bm{y}$ and used for solving:
@@
Figure 3: Gaussian noisy image and its parameter maps.
@@ II-A Image patch dataset & network architecture
As a training dataset, we considered $300$ images extracted from the popular Berkeley greyscale image dataset (BSD300), covering a diverse range of subjects (from faces, to natural landscapes etc). We preliminary split the dataset into 200/50/50 images for training/validation and testing. Overlapping patches of size $32\times 32$ were extracted, with a sliding step set to 16 pixels, either to the right or downward, thus letting the patches to overlap by 25% or 50%. From each extracted patch, several noisy patches generated as different realisations of fixed Gaussian/Poisson noise distributions were generated. To create labels, i.e. reference values $\tilde{\mu}_{i}$ to be used for training, problem (7) was solved repeatedly (see Section II-B). To facilitate the learning process, the outlier values that did not fall in the range $(0,1.5\times\text{IQR}]$ were removed. The overall samples in the dataset for Poisson denoising are 144,870, for Gaussian denoising are 165,300. Two separate neural networks were trained for Gaussian and Poisson denoising. Both architectures have approximately $2.8\times 10^{6}$ model parameters, as detailed in Table I. As a training loss the mean squared error (MSE) between the predicted $\mu$ and the reference $\tilde{\mu}$ was used. Adam optimizer was used with an initial learning rate of $10^{-5}$ and an exponential decay rate of factor $\gamma=0.8$, along with a weight decay of parameter $10^{-4}$. Training and validation were performed over 30 epochs with a batch size of 256.
@@ II-C Noise model selection: architecture
For choosing the appropriate data term (and, consequently, the reconstruction pipeline/algorithm) for computing labels, a binary noise classification task (Gaussian VS. Poisson) was performed. For that, patches of size $64\times 64$ were used. In comparison to the denoising setting, larger patch size allows here for richer contextual information for distinguishing between noise types. For training the classification network, each clean patch of the dataset was injected with noise of Gaussian/Poisson distribution with varying intensity levels with a total number of 37,800 samples. Specifically, noise intensities were set to $\sigma^{2}=\{0.01,0.02,0.03\}$ for Gaussian noise and $\alpha=\{45,30,10\}$ for Poisson noise.
The classification network employed was trained by minimising the cross-entropy loss between predicted and ground-truth classes. The total number of parameters employed is approximately $1.7\times 10^{7}$. As far as the optimization is concerned, the same setting as the one described for training the parameter estimation network was employed.
@@ III Numerical results
The proposed approach is thus divided into two sequential steps: first the noise classifier network is applied to compute whether Gaussian ($\delta=1$) or Poisson ($\delta=0$) noise is observed in the image at hand. Then such parameter is plugged into (8), so that the trained network $f_{\bm{\theta}}$ computing pixel-wise the parameter map $\bm{\mu}$ is applied in a sliding-window fashion. In the following we report the results obtained in correspondence of a Gaussian denoising task with fixed variance $\sigma^{2}=0.01$ and a Poisson denoising task with fixed parameter $\alpha=30$ after noise model selection is performed.
@@ III-A Noise model selection: performance
the results in Table II, covering a range from low noise to higher noise levels.
@@ III-B Parameter map estimation: performance
To validate the performance of the proposed approach, comparisons were performed on a set of test images between the space-adaptive (8) and the scalar approach $\mu_{i}\equiv\mu$ in terms of computational times, peak signal noise ratio (PSNR) and the structural similarity index measure (SSIM) quality metrics.
@@
We report in the following some results obtained in correspondence of a noisy image $\bm{y}$ corrupted with Gaussian noise with zero mean and variance $\sigma^{2}=0.01$ and show improvements w.r.t. the space-variant approach, see Fig. (4). Improvements upon SSIM and PSNR metrics are visible compared to the scalar approach, showing, in particular texture enhancement and stronger smoothing in homogeneous areas. We tested the approach over a dataset of $50$ validation images, observing an average SSIM improvement of 0.017 and PSNR improvement of 0.35 w.r.t. the scalar approach. In terms of computational times, inference required on average $\sim$ 187 seconds of CPU time, in comparison of $\sim$ 183 seconds required for computing the solution of the scalar approach with, alongside, optimal scalar parameter selection via golden section which requires direct comparison with the ground truth image and is unfeasible in case of the proposed pixel-dependent parameter map computation.

### v2

@@ Abstract
We consider a patch-based learning approach defined in terms of neural networks to estimate spatially adaptive regularisation parameter maps for image denoising with weighted Total Variation (TV) and test it to situations when the noise distribution is unknown. As an example, we consider situations where noise could be either Gaussian or Poisson and perform preliminary model selection by a standard binary classification network. Then, we define a patch-based approach where at each image pixel an optimal weighting between TV regularisation and the corresponding data fidelity is learned in a supervised way using reference natural image patches upon optimisation of SSIM and in a sliding window fashion. Extensive numerical results are reported for both noise models, showing significant improvement w.r.t. results obtained by means of optimal scalar regularisation.
@@ I Introduction
Traditionally, the parameter $\mu$ is a scalar value, so that regularisation is applied uniformly across the entire image domain. More recent approaches (see, e.g., [7]) consider weighted, that is spatially-adaptive, variants of TV-based models where a pixel-dependent weighting between regularisation and data term is used to account for differences between textured and geometric features, thus avoiding unbalanced effects corresponding to a global selection of $\mu$, see Fig. 1, where the regularisation parameter maps is computed in a patch-wise fashion and optimal weighting values are computed on an exemplar image corrupted by Gaussian noise with $\sigma^{2}=0.01$.
@@ I-A Contribution
We consider a patch-based learning approach for estimating adaptive, i.e. spatially-varying regularisation maps for the problem of image denoising with unknown noise distribution. The regularisation map $\bm{\mu}_{\text{map}}$ is estimated in a supervised learning fashion where, for each noisy patch $\bm{y}_{p}$, the map $\bm{y}_{p}\mapsto\mu_{i}$ providing an estimate of the regularisation parameter to-be-located in the pixel $i=1,\ldots,n$ placed at the centre of $\bm{y}_{p}$ is parametrised by a neural network $f_{\theta}$ which, at inference time, operates in a sliding-window fashion, taking as inputs patches of fixed size. An example of the proposed map estimation can be appreciated in Fig. 2.
Noise model selection is performed using a standard binary classification network $g_{\psi}$ able to distinguish between Gaussian and Poisson noise. Several numerical results are reported showing visual and statistical improvements w.r.t. non-adaptive regularisation models.
@@ II Patch-based learning of regularisation maps
We consider in the following an adaptive approach that assigns to each pixel $i\in\left\{1,\ldots n\right\}$ of a noisy image $\bm{y}$ corrupted by Gaussian/Poisson noise a regularisation parameter $\mu_{i}$ computed by means of a suitable training of a deep neural network leveraging information on a small patch centered in the pixel itself. More precisely, given a patch $\bm{y}_{p_{i}}$ extracted by $\bm{y}$ around $i\in\left\{1,\ldots n\right\}$, we consider the following patch-wise version of problem (3) suited to deal with Gaussian/Poisson which reads:
@@
where $\delta\in\left\{0,1\right\}$ identifies (upon suitable training) whether Gaussian/Poisson noise is observed in the image, so that a tailored data term is used for reconstruction. The parameter $\mu_{i}$ is the one we would like to estimate to construct the desired regularisation map. We now parametrise the map $\bm{y}_{p_{i}}\mapsto\mu_{i}$ using a deep neural network $f_{\bm{\theta}}$. In the following, we describe the supervised-learning strategy which, given a dataset of clean/noisy natural image patches $\left\{(\tilde{\bm{x}}_{p}^{k},\bm{y}^{k}_{p})\right\}_{k=1}^{K}$ learn such map. At inference time, note that this map will be applied in a sliding-window fashion, see Fig. 2, to an unobserved image $\bm{y}$ and used for solving:
@@ II-A Image patch dataset & network architecture
As a training dataset, we considered $300$ images extracted from the popular Berkeley greyscale image dataset (BSD300), covering a diverse range of subjects (from faces, to natural landscapes etc). We preliminary split the dataset into 200/50/50 images for training/validation and testing. Overlapping patches of size $32\times 32$ were extracted, with a sliding step set to 16 pixels, either to the right or downward, thus letting the patches to overlap by 25% or 50%. From each extracted patch, several noisy patches generated as different realisations of fixed Gaussian/Poisson noise distributions were generated. To create labels, i.e. reference values $\tilde{\mu}_{i}$ to be used for training, problem (7) was solved repeatedly (see Section II-B). To facilitate the learning process, the outlier values that did not fall in the range $(0,1.5\times\text{IQR}]$ were removed. The overall samples in the dataset for Poisson denoising are 144,870, for Gaussian denoising are 165,300. Two separate neural networks were trained for Gaussian and Poisson denoising. Both architectures have approximately $2.8\times 10^{6}$ model parameters, as detailed in Table I. Its layer architecture is based on the one of the well-known AlexNet, with modifications aimed at reducing the total number of parameters, so as to mitigate overfitting (via dropout layers). As a training loss the mean squared error (MSE) between the predicted $\mu$ and the reference $\tilde{\mu}$ was used. Adam optimizer was used with an initial learning rate of $10^{-5}$ and an exponential decay rate of factor $\gamma=0.8$, along with a weight decay of parameter $10^{-4}$. Training and validation were performed over 30 epochs with a batch size of 256.
@@ II-C Noise model selection: architecture
For choosing the appropriate data term (and, consequently, the reconstruction pipeline/algorithm) for computing labels, a binary noise classification task (Gaussian VS. Poisson) was performed. For that, patches of size $64\times 64$ were used. In comparison to the denoising setting, larger patch size allows here for richer contextual information for distinguishing between noise types. For training the classification network, each clean patch of the dataset was injected with noise of Gaussian/Poisson distribution with varying intensity levels with a total number of 37,800 samples. Specifically, noise intensities were set to $\sigma^{2}=\{0.01,0.02,0.03\}$ for Gaussian noise and $\alpha=\{45,30,10\}$ for Poisson noise. While structurally similar to the architecture detailed in Table I, the classification network differs in some key aspects: it applies higher channel counts in earlier layers (i.e. bigger patches), it lacks a MaxPool2D layer and it ends with a softmax output layer for binary classification. The increased number of channels thus results in a greater number of units in the dense layers so that the total number of parameters employed is approximately $1.7\times 10^{7}$. Training was performed by minimising the cross-entropy loss between predicted and ground-truth classes. As far as the optimization is concerned, the same setting as the one described for training the parameter estimation network was employed.
@@ III Numerical results
The proposed approach is thus divided into two sequential steps: first the noise classifier network is applied to compute whether Gaussian ($\delta=1$) or Poisson ($\delta=0$) noise is observed in the image at hand. Then such parameter is plugged into (8), so that the trained network $f_{\bm{\theta}}$ computing pixel-wise the parameter map $\bm{\mu}$ is applied in a sliding-window fashion. A comprehensive overview of our proposed pipeline is shown in Fig. 2. In the following we report the results obtained in correspondence of a Gaussian denoising task with fixed variance $\sigma^{2}=0.01$ and a Poisson denoising task with fixed parameter $\alpha=30$ after noise model selection is performed.
Figure 2: The input noisy image is taken as input by a noise classifier to identify the noise type. Based on the classification result, the corresponding denoising routine is used to estimate a pixel-wise regularisation map in a sliding-window fashion.
@@ III-B Parameter map estimation: performance
To validate the performance of the proposed approach, comparisons were performed on a set of test images between the space-variant method (8) exploiting locally adaptive regularisation, and a scalar approach where a single parameter $\mu_{i}\equiv\mu$ is computed—via SSIM maximisation (see Sec. 2B)—on the entire image. The two approaches are compared in terms of computational time, PSNR and SSIM.
@@
We report in the following some results obtained in correspondence of a noisy image $\bm{y}$ corrupted with Gaussian noise with zero mean and variance $\sigma^{2}=0.01$ and show improvements w.r.t. the scalar approach, see Fig. (3). Improvements upon SSIM and PSNR metrics are visible compared to the scalar approach, showing, in particular texture enhancement and stronger smoothing in homogeneous areas. We tested the approach over a dataset of $50$ validation images, observing an average SSIM improvement of 0.017 and PSNR improvement of 0.35 w.r.t. the scalar approach. In terms of computational times, inference required on average $\sim$ 187 seconds of computational time, in comparison of $\sim$ 183 seconds required for computing the solution of the scalar approach with, alongside, optimal scalar parameter selection via golden section which requires direct comparison with the ground truth image and is unfeasible in case of the proposed pixel-dependent parameter map computation.